A world-class sprinter can burst out of the blocks to essentially top speed (of about 11.5 m/s) in the first 15.0 m of the race. What is the average acceleration of this sprinter, and how long does it take her to reach that speed?


1
Expert's answer
2021-10-08T09:22:41-0400

Given:

"v_i=0\:\rm m/s"

"v_f=11.5\:\rm m/s"

"d=15.0\:\rm m"


The distance traveled by a car is given by


"d=\frac{v_f^2-v_i^2}{2a}=\frac{v_f+v_i}{2}t"

Hence, the acceleration of a car

"a=\frac{v_f^2-v_i^2}{2d}=\rm\frac{11.5^2-0^2}{2*15.0}=4.41\:\rm m/s^2"

The time of motion

"t=\frac{2d}{v_f+v_i}=\rm\frac{2*15.0}{11.5+0}=2.61\:\rm s"